OpenStudy (anonymous):

x+3y=9 and 4x-2y=8 how to this slope y intercept and x intercept

OpenStudy (anonymous):

What's your question?

OpenStudy (anonymous):

Try changing the equations to be the formula Y=mx+b

OpenStudy (anonymous):

X+3y=9 Change it to y=mx+b Subtract x from both sides 3y=-x+9 Answers would be: Y=-1/3x+3 Y- intercept is 3 Slope is -1/3